# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ V-B Localization and Tracking
Existing localization and tracking solutions often require continuous communication and processing, leading to significant energy consumption, particularly in battery-constrained devices [219]. Additionally, it faces challenges in scaling to support a large number of devices simultaneously, leading to increased computational complexity and resource consumption. QC facilitates simultaneous processing of multiple possibilities, which accelerates and enhances localization and tracking tasks. This capability finds applications in various scenarios, including tracking mobile user trajectories [220], managing wireless sensor networks [221, 222], and coordinating multiple UAV networks [223]. User trajectory prediction in [220] involved two components: reservoir computing and QC. Compared to training simple recurrent neural networks, reservoir computing was computationally more efficient because only the output layer weights were trainable. However, achieving optimal performance in reservoir computing requires careful selection of reservoir weights to create complex and nonlinear dynamics. QC was employed to generate such complex reservoir dynamics, mapping input time series into a higher-dimensional computational space consisting of dynamical states. Once high-dimensional dynamical states were obtained, a simple linear regression was applied to train the output weights, enabling efficient prediction of mobile users’ trajectories.
@@
QC also finds extensive applications in localization and tracking within sensor networks, particularly addressing scalability challenges. The approach in [221] enhanced the differential crossover in QPSO to address nonlinear localization problems influenced by node density and coverage in IoT. By utilizing the differential evolution operator, it mitigated group movements within narrow ranges, thus preventing convergence to local optima and enhancing global search capabilities. The cross operator facilitated information exchange among individuals within a group, while exceptional genes were moderately retained to sustain the evolutionary process, thereby incorporating both proactive and reactive features. In contrast to traditional location scenarios as presented in [221], [222] addressed the localization with multiple mobile robots equipped with sensors aiming to locate a specific source node. By modeling each robot as a double integrator, [222] converted the source location challenge into a multi-robot path planning task, considering constraints such as collisions between robots and obstacles, as well as collisions among robots. A framework for the mobile policy of the robots was devised. And an adaptive weight strategy, a swarm timely update mechanism, and a leading-following behavior were integrated into QPSO. These enhancements aimed to enable multiple robots to adhere to the mobile policy effectively, ensuring robustness and efficiency in their search.
@@
QC, paired with AI, presents new ways to tackle location and tracking tasks in UAV networks [223]. In [223], a DRL approach was proposed for controlling vertical take-off and landing aircraft in the presence of wind disturbances. The tracking control problem was framed as MDP. To enhance control accuracy despite wind disturbances, various wind fields were incorporated into the learning environment to broaden exploration and mimic the impact of winds on flight control. To ensure both tracking accuracy and practical implementation, a quantum-inspired experience replay strategy was introduced based on QC principles. This strategy featured a preparation operation scheme to foster exploration and accelerate convergence, along with a depreciation operation method to enhance sample diversity and bolster system robustness.
@@
Previous efforts to locate and track devices relied on sensors or GPS, which was practical for scenarios like indoor person tracking with only a cellphone. Existing solutions for such an application typically utilize received signal strength (RSS) to analyze the location, but faces a challenge of scalability to worldwide and algorithm complexity. The work [224] provided a fast matching quantum fingerprinting algorithm by calculating the quantum cosine similarity between online RSS vector at an unknown location and each RSS vector in the cellular fingerprint. Its efficiency was validated in the testbed IBM Quantum Machine Simulator and cellular BSs. While it is an innovative approach, the matching accuracy can be further enhanced by utilizing channel state information (CSI) instead of RSS. However, acquiring CSI presents a practical challenge due to hardware limitations. Moreover, RSS, while commonly used, is susceptible to noise factors like pedestrian interference, which is not accounted for in the solution.
@@ V-C Video Streaming
Video serves as a primary means of conveying information through visual and auditory media and finds widespread applications in entertainment, education, and information dissemination. A typical video system includes several components such as video encoding/decoding [225], network transmission [226], and buffer [227], etc. differing significantly from conventional data transmission. Additionally, emerging technologies such as multi-view videos [228], free-viewpoint videos [229], VR/360 videos [230], and point cloud videos [231, 232] have surfaced in recent years. These innovations allow users to choose the viewing angle, enabling them to selectively watch specific parts of the video during the viewing process, for example, in the case of 360 videos/point cloud videos, where users can switch perspectives over time.
In wireless communications and networking, one classical issue is adaptive bitrate (ABR) control in order to ensure the video streaming quality. The work [233] maximized the user QoE by increasing the average bitrate and decreasing the video rebuffering events. Specifically, the proposal formulated the control problem into a QUBO model from three aspects: video quality, bitrate change, and buffer condition, which was solved in the platform Digital Annealer. Representative applications, VR and Metaverse in 6G and beyond, also employed QC algorithms to improve performance. Ultrahigh viewport rendering demands and excessive terminal energy consumption restrict the application of wireless VR. The work [234] considered content correlation between VR equipments, fluctuating channel conditions, and VR QoE, by formulating the joint problem of the viewport rendering offloading, computing, and spectrum resource allocation. The formulation was solved by proposed RL algorithm and quantum parallelism was integrated into the RL to overcome the problem of high data dimension.
@@ VI-A QC in Wireless Communications and Networking with Multiple Variables
QC has been applied in wireless communications and networking, primarily focusing on individual variables. However, tackling optimization problems involving multiple variables poses a more formidable challenge. Addressing multiple variables in wireless systems presents significant hurdles for optimization and control. The exponential increase in dimensionality of optimization problems renders them computationally intensive and impractical for real-time decision-making. Balancing conflicting objectives, such as maximizing throughput while minimizing latency and interference, becomes intricate. Moreover, dynamic environments exacerbate the complexity, introducing uncertainty in predicting optimal configurations. Complex interactions among variables, compounded by resource constraints, further compound the challenge. Efficient algorithms that integrate optimization theory, signal processing, and machine learning are essential to effectively tackle these challenges.
On the other hand, AI-driven quantum algorithms sometimes require datasets for training, which is currently hindered by insufficient available datasets.
@@ VI-B QC in Wireless Communications and Networking for Particular Services
Introducing QC into wireless communications and networking for specific services poses several challenges. For instance, services like video streaming demand stringent requirements beyond conventional wireless systems, such as low latency, high throughput, and reliable transmission. Optimizing these parameters using quantum algorithms becomes more challenging due to the complexity introduced by the additional service requirements. Quantum algorithms must navigate the trade-offs between different objectives, such as maximizing video quality while minimizing latency and packet loss. Furthermore, the dynamic nature of video traffic adds an extra layer of complexity, requiring adaptive optimization strategies that can efficiently allocate resources in real-time. Additionally, ensuring the security and privacy of video data transmission becomes more critical with the integration of QC, necessitating robust encryption schemes and secure communication protocols that can withstand quantum attacks. Overall, integrating QC into wireless communications and networking or specific services like video streaming presents challenges in balancing competing objectives, adapting to dynamic traffic patterns, and ensuring robust security measures, requiring innovative solutions tailored to the unique requirements of each service.
@@ VI-D QC Cloud Services in Wireless Communications and Networking
As QC becomes increasingly available to the public through cloud services, there is a growing need to ensure these services can be effectively supported in wireless networks in the near future [235]. Factors such as communication latency and stability will be critical considerations in meeting network requirements. Furthermore, the integration of QC into the classical computing pipeline will involve determining which components are most suitable for quantum processing to maximize benefits. This includes addressing the overhead associated with transitioning between classical and QC, a crucial aspect for optimizing overall efficiency and performance.
@@ VII Conclusions
The advent of QC technology signals a significant revolution in wireless communications and networking. To facilitate the exploration of its applications in wireless communications and networking, we provided a tutorial on QC, covering its basics, characteristics, and popular QC algorithms. Additionally, we introduced the applications of QC in communication and networking, followed by its applications in miscellaneous areas such as security and privacy, localization and tracking, and video streaming. We believe this is a very promising direction and have discussed the remaining open issues.

### v3

@@ V-B Localization and Tracking
Existing localization and tracking solutions often require continuous communication and processing, leading to significant energy consumption, particularly in battery-constrained devices [197]. Additionally, it faces challenges in scaling to support a large number of devices simultaneously, leading to increased computational complexity and resource consumption. QC facilitates simultaneous processing of multiple possibilities, which accelerates and enhances localization and tracking tasks. This capability finds applications in various scenarios, including tracking mobile user trajectories [198], managing wireless sensor networks [199, 200], and coordinating multiple UAV networks [201]. User trajectory prediction in [198] involves two components: reservoir computing and QC. Compared to training simple recurrent neural networks, reservoir computing is computationally more efficient because only the output layer weights are trainable. However, achieving optimal performance in reservoir computing requires careful selection of reservoir weights to create complex and nonlinear dynamics. QC is employed to generate such complex reservoir dynamics, mapping input time series into a higher-dimensional computational space consisting of dynamical states. Once high-dimensional dynamical states are obtained, a simple linear regression is applied to train the output weights, enabling efficient prediction of mobile users’ trajectories.
@@
QC also finds extensive applications in localization and tracking within sensor networks, particularly addressing scalability challenges. The approach in [199] enhances the differential crossover in QPSO to address nonlinear localization problems influenced by node density and coverage in IoT. By utilizing the differential evolution operator, it mitigates group movements within narrow ranges, thus preventing convergence to local optima and enhancing global search capabilities. The cross operator facilitates information exchange among individuals within a group, while exceptional genes are moderately retained to sustain the evolutionary process, thereby incorporating both proactive and reactive features. In contrast to traditional location scenarios as presented in [199], [200] addresses the localization with multiple mobile robots equipped with sensors aiming to locate a specific source node. By modeling each robot as a double integrator, [200] converts the source location challenge into a multi-robot path planning task, considering constraints such as collisions between robots and obstacles, as well as collisions among robots. A framework for the mobile policy of the robots is devised. And an adaptive weight strategy, a swarm timely update mechanism, and a leading-following behavior are integrated into QPSO. These enhancements aims to enable multiple robots to adhere to the mobile policy effectively, ensuring robustness and efficiency in their search.
@@
QC, paired with AI, presents new ways to tackle location and tracking tasks in UAV networks [201]. In [201], a DRL approach is proposed for controlling vertical take-off and landing aircraft in the presence of wind disturbances. The tracking control problem is framed as MDP. To enhance control accuracy despite wind disturbances, various wind fields are incorporated into the learning environment to broaden exploration and mimic the impact of winds on flight control. To ensure both tracking accuracy and practical implementation, a quantum-inspired experience replay strategy is introduced based on QC principles. This strategy features a preparation operation scheme to foster exploration and accelerate convergence, along with a depreciation operation method to enhance sample diversity and bolster system robustness.
@@
Previous efforts to locate and track devices relies on sensors or GPS, which is impractical for scenarios like indoor person tracking with only a cellphone. Existing solutions for such an application typically utilize received signal strength (RSS) to analyze the location, but face challenges of scalability to worldwide and algorithm complexity. The work [202] provides a fast matching quantum fingerprinting algorithm by calculating the quantum cosine similarity between online RSS vector at an unknown location and each RSS vector in a cellular fingerprint. Its efficiency is validated in the testbed IBM quantum machine simulator and cellular BSs. While it is an innovative approach, the matching accuracy can be further enhanced by utilizing channel state information (CSI) instead of RSS. However, acquiring CSI presents a practical challenge due to hardware limitations. Moreover, RSS, while widely used, is susceptible to noise factors like pedestrian interference, which is not accounted for in the solution.
@@ V-C Video Streaming
Video serves as a primary means of conveying information through visual and auditory media and finds widespread applications in entertainment, education, and information dissemination. A typical video system includes several components such as video encoding/decoding, network transmission, and buffer, differing significantly from conventional data transmission. Additionally, emerging technologies [203, 204] such as multi-view videos, free-viewpoint videos, VR/360 videos, and point cloud videos, have surfaced in recent years. These innovations allow users to choose viewing angles, enabling them to selectively watch specific video parts during viewing processes, for example, in the case of 360 videos/point cloud videos, where users can switch perspectives over time.
In wireless communications and networking, one classical issue is adaptive bitrate control in order to ensure video streaming quality. The work [205] maximizes user QoE by increasing average bitrate and decreasing video rebuffering events. Specifically, the proposal formulates the control problem into a QUBO model from three aspects: video quality, bitrate change, and buffer condition, which is solved in the platform Digital Annealer. Representative applications, VR and Metaverse in 6G and beyond, also employ QC algorithms to improve performance. Ultrahigh viewport rendering demands and excessive terminal energy consumption restrict the application of wireless VR. The work [206] considers content correlation between VR equipments, fluctuating channel conditions, and VR QoE, by formulating the joint problem of viewport rendering offloading, computing, and spectrum resource allocation. The formulation is solved by proposed RL algorithm and quantum parallelism is integrated into the RL to overcome the problem of high data dimension.
@@ VI-A QC in Wireless Communications and Networking with Multiple Variables
QC has been applied in wireless communications and networking, primarily focusing on individual variables. However, tackling optimization problems involving multiple variables poses a more formidable challenge while considering network characteristics, such as high dimensionality, conflicting objectives, resource constraints and so forth. Addressing multiple variables in wireless systems presents significant hurdles for optimization and control. The exponential increase in dimensionality of optimization problems renders them computationally intensive and impractical for real-time decision-making. Balancing conflicting objectives, such as maximizing throughput while minimizing latency and interference, becomes intricate. Moreover, dynamic environments exacerbate the complexity, introducing uncertainty in predicting optimal configurations. Complex interactions among variables, compounded by resource constraints, further compound the challenge. Efficient algorithms that integrate optimization theory, signal processing, and ML is essential to effectively tackle these challenges. On the other hand, AI-driven quantum algorithms sometimes require datasets for training, which is currently hindered by insufficient available datasets.
The challenge of handling multiple variables in wireless networks using QC can be addressed by integrating hybrid quantum-classical optimization frameworks. These frameworks utilize classical techniques such as ML and optimization theory to manage simpler or well-understood parts of the problem, while quantum algorithms like QAOA focus on the most computationally intensive tasks. Additionally, techniques like dimensionality reduction and adaptive quantum algorithms can help mitigate the exponential growth of variables and provide real-time solutions. Developing quantum-inspired algorithms that efficiently handle conflicting objectives, such as balancing throughput, latency, and interference, will also be critical for the challenge.
@@ VI-B QC in Wireless Communications and Networking for Particular Services
Introducing QC into wireless communications and networking for specific services poses several challenges while considering stringent service requirements, complex trade-offs, dynamic traffic patterns, and so forth. For instance, services like video streaming demand stringent requirements beyond conventional wireless systems, such as low latency, high throughput, and reliable transmission. Optimizing these parameters using quantum algorithms becomes more challenging due to the complexity introduced by the additional service requirements. Quantum algorithms must navigate the trade-offs between different objectives, such as maximizing video quality while minimizing latency and packet loss. Furthermore, the dynamic nature of video traffic adds an extra layer of complexity, requiring adaptive optimization strategies that can efficiently allocate resources in real-time. Additionally, ensuring the security and privacy of video data transmission becomes more critical with the integration of QC, necessitating robust encryption schemes and secure communication protocols that can withstand quantum attacks. Overall, integrating QC into wireless communications and networking or specific services like video streaming presents challenges in balancing competing objectives, adapting to dynamic traffic patterns, and ensuring robust security measures, requiring innovative solutions tailored to the unique requirements of each service.
To optimize QC in wireless networks for specific services like video streaming, hybrid quantum-classical algorithms can be developed to address the trade-offs between quality, latency, and packet loss. QML can predict traffic patterns, allowing for adaptive resource allocation in real-time. To ensure robust security in these systems, quantum-resistant encryption protocols must be integrated to safeguard data from quantum attacks. Additionally, early-stage quantum simulations can be used to test and refine these solutions, paving the way for practical implementations in complex service scenarios.
@@ VI-C Understanding and Utilizing QC in Communications and Networking
Addressing the complexity and understanding of QC in wireless communications requires the creation of educational tools and resources that break down quantum concepts into more accessible terms for professionals in the field. By developing user-friendly interfaces and software that abstract complex quantum operations, wireless engineers can more easily experiment with and adopt QC techniques. Collaborative efforts between academia and industry, along with standardized training programs, will help bridge the knowledge gap and accelerate the practical use of QC in real-world wireless systems.
@@ VI-D QC Cloud Services in Wireless Communications and Networking
As QC becomes increasingly available to the public through cloud services, there is a growing need to ensure these services can be effectively supported in wireless networks in the near future. Factors such as communication latency and stability will be critical considerations in meeting network requirements. Furthermore, the integration of QC into the classical computing pipeline will involve determining which components are most suitable for quantum processing to maximize benefits. This includes addressing the overhead associated with transitioning between classical and QC, a crucial aspect for optimizing overall efficiency and performance.
To support QC cloud services over wireless networks, a focus on optimizing communication latency and stability is key. One potential solution is to deploy quantum edge computing to localize quantum processing tasks closer to users, reducing network overhead and latency. Additionally, integrating QC with classical systems should prioritize tasks most suited to quantum processing, thereby maximizing performance while minimizing the overhead associated with switching between classical and quantum systems. Improved network protocols that efficiently manage these transitions will be necessary for enhancing the performance of QC cloud services in wireless networks.
VI-E Integration of DRL with QC
While many studies have explored the integration of DRL with QC to address issues in wireless networks, expanding on the challenges that QC faces in DRL requires an exploration of the key areas where these two fields intersect. 1) Measurement and information loss: In QC, a quantum state collapses upon measurement, leading to a loss of information. Since DRL relies heavily on learning from observed states and rewards, this presents a challenge when attempting to optimize quantum systems in real-time without collapsing the states. 2) Noise and decoherence: Quantum systems are highly sensitive to environmental noise and decoherence, which degrades the quality of quantum information and calculations. DRL algorithms that attempt to optimize quantum circuits may struggle with learning stable policies in the presence of such noise. 3) Quantum reward structures: DRL relies on reward signals to guide the learning process. In quantum systems, the nature of the reward structures may be less intuitive than classical systems due to the probabilistic nature of quantum operations, leading to challenges in defining meaningful rewards.
To overcome the challenges of integrating DRL with QC in wireless networks, hybrid frameworks can be designed that separate quantum and classical tasks. Classical DRL algorithms can handle observation and learning, while quantum algorithms focus on optimizing decision-making processes in high-dimensional spaces. Developing quantum error correction techniques and robust algorithms that account for quantum noise and decoherence will also help stabilize learning. Quantum-inspired reward structures that align with DRL principles must be created to enhance the learning process while ensuring the stability and reliability of quantum systems.
@@ VII Conclusions
This paper presents a comprehensive survey of QC in wireless communications and networking, providing an in-depth overview of the field’s fundamentals, advancements, and application potentials. By integrating QC into the optimization processes of wireless networks, significant improvements can be made in areas such as resource allocation, power management, edge computing, and security.
The primary contributions of this work lie in its interdisciplinary coverage, bridging quantum mechanics and wireless communications, and in providing a clear roadmap for applying QC algorithms to various aspects of wireless networking. The survey not only highlights the theoretical frameworks of QC but also examines its practical implications in real-world scenarios, such as edge computing and content caching. Additionally, it identifies key challenges and potential solutions for leveraging quantum algorithms in next-generation wireless networks.
The ongoing development of QC platforms, along with their application in wireless communications, opens up promising avenues for future research. Specific areas such as QRL, edge intelligence with QC, and advanced security protocols powered by quantum technologies will likely evolve further. Future research should focus on minimizing hardware noise and gate errors to improve algorithmic efficiency, as well as exploring hybrid classical-quantum frameworks for large-scale implementations. Expected outcomes from these advancements include faster optimization processes, more robust security mechanisms, and enhanced communication systems capable of meeting the demands of the upcoming 6G era.